Zelda borrowed R5760,00

 from a bank for 21 months. For the discount loan she only received R4752,00

. Determine: 


    (i)   the yearly discount rate, d


    (ii)   the equivalent yearly simple interest rate, i



Give the answers as percentages and round them to two decimal places.


1
Expert's answer
2021-10-12T13:44:09-0400

In this we have to calculate present value on based on discount rate.


i

Discounted value = Present value=475200

Borrowed amount =576000

Period = 21 months =1.75 years 

r = Annual rate of interest

"Present\space value = \frac{Future\space value}{(1+r)^{1.75}}"

"475200=\frac{576000}{(1+r)^{1.75}}"

"(1+r)^{1.75}=1.2121212121"

r=11.62%

This is annual rate of interest.


ii

"Simple\space discount rate = \frac{Discount}{ borrowed\space amount }\times\frac{ 12\space months}{ Loan \space period }"


Discount =576000-475200=$100800


Simple interest rate "=\frac{101800}{576000} \times \frac{12}{21}=10\%"


Simple interest rate=10%